A principal of $5000 was invested in a savings account for 4 years. If the interest earned for the period was $400, what was the interest rate?
Found 2 solutions by stanbon, tutorcecilia:
Answer by stanbon(75887)

A principal of $5000 was invested in a savings account for 4 years. If the interest earned for the period was $400, what was the interest rate?
Interest = Principal * rate * time
400 = 5000*r*4
1/50=r
r=2/100
r=2%
Cheers,
Stan H.

Answer by tutorcecilia(2152)

I=Prt [use the formula for computing simple interest]
$400=(5000)(r)(4) [solve for r]
400=20,000r
400/20,000=r
.02=r
.02x100%=2%